# Sectorial covers over fanifolds

### v2

@@ 3.1. Fanifolds
    $\Phi$ is conical in the complement of a compact subset.
@@
One can identify the exit path category $\operatorname{Exit}(\Sigma)$ with $\operatorname{Fan}^{\twoheadrightarrow}_{\Sigma/}$ as a poset via
$\sigma\mapsto[\Sigma\mapsto\Sigma/\sigma].$
In addition, the normal geometry to $\sigma$ is the geometry of $\Sigma/\sigma$. Due to the condition (v), $\operatorname{Exit}_{S}(\Phi)$ is equivalent to the poset $\operatorname{Exit}(\Sigma_{S})$, the full subcategory of exit paths starting at $S$ contained inside a sufficiently small neighborhood of $S$.
@@
Suppose that $\Phi_{k-1}$ has the desired filtration $\Phi_{0}\subset\Phi_{1}\subset\cdots\subset\Phi_{k-1}$, where $\Phi_{0}$ is the disjoint union $\bigsqcup_{P}\Sigma_{P}$ for all 00-strata $P\subset\Phi$ equipped with the canonically induced fanifold structure. The ideal boundary $\partial_{\infty}S$ might have some subset $\partial_{in}S$ which is in the direction to the interior of $\Phi$. Perform the gluing $\Phi_{k-1}\#_{(\Sigma_{S}\times\partial_{in}S)}(\Sigma_{S}\times S)$ which equals $\Sigma_{S}\times S$ when $\Phi_{k-1}$ is empty and equals $\Phi_{k-1}$ unless $S$ is an interior $k$-stratum, i.e., $\partial_{in}S=\partial S_{\circ}$ where $S_{\circ}=S\setminus\Phi_{k-1}$ is a manifold-with-boundary. Let
@@ 3.2. Weinstein handle attachments
where the Liouville flow on $W$ gets identified with the translation action on $\mathbb{R}_{\leq 0}$. By [Eli, Proposition 2.9] one can modify the Weinstein structure near $\mathcal{L}$ so that the Liouville flow gets identified with the cotangent scaling on $T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$. We denote by $\widetilde{W}$ its conic completion. Then $\eta$ induces a neiborhood $\widetilde{\eta}\colon\operatorname{Nbd}_{\partial_{\infty}\widetilde{W}}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}$ which extends to a neighborhood $\widetilde{\xi}\colon\operatorname{Nbd}_{\widetilde{W}}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}$.
Remark 3.6.
From [GPS1, Lemma 2.13] it follows that $\widetilde{W}$ is a Weinstein sector.
@@
where $\operatorname{Core}(W,\mathcal{L})$ is the relative skeleton of $(W,\lambda)$ associated with $\mathcal{L}$, i.e., the disjoint union $\operatorname{Core}(W)\sqcup\mathbb{R}\mathcal{L}\subset W$ of $\operatorname{Core}(W)$ and the saturation of $\mathcal{L}$ by the Liouville flow.
@@
Let $W$ be a Weinstein domain with a smooth Legendrian $\mathcal{L}\subset\partial_{\infty}W$. Then a Lagrangian $\mathbb{L}\subset W$ is biconic along $\mathcal{L}$ if and only if it is conic and
$\xi(\mathbb{L}\cap\operatorname{Nbd}_{W}(\mathcal{L}))\subset T^{*}\mathcal{L}\times\{0\}\times\mathbb{R}_{\leq 0}\subset J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}\cong T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$
for some $\xi\colon\operatorname{Nbd}_{W}(\mathcal{L})\hookrightarrow T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$, where the Liouville flow on $W$ gets identified with the translation action on $\mathbb{R}_{\leq 0}$.
@@
For a closed manifold $\widehat{M}$ and a compact manifold-with-boundary $S$, consider the Weinstein manifold-with-boundary $\widetilde{W^{\prime}}=T^{*}\widehat{M}\times T^{*}S$ with a smooth Legendrian $\mathcal{L}=\widehat{M}\times\partial S$ taken to be a subset of the zero section. Here, we equip $T^{*}S$ the tautological Liouville structure so that the Liouville vector field is the generator of fiberwise radial dilation. Then the Liouville flow on $\widetilde{W^{\prime}}$ near $\mathcal{L}$ is the cotangent scaling. We write $W^{\prime}$ for a Weinstein domain completing to $\widetilde{W^{\prime}}$. One can check that the above gluing procedure carries over, although $\mathcal{L}$ does not belong to $\partial_{\infty}W^{\prime}$.
@@
Definition 3.10 ([FLTZ12, Section 3.1]).
Let $\Sigma\subset M_{\mathbb{R}}$ be a rational polyhedral fan and $\widehat{M}$ the real $n$-torus $\operatorname{Hom}(M,\mathbb{R}/\mathbb{Z})=M^{\vee}_{\mathbb{R}}/M^{\vee}$. The FLTZ Lagrangian is the union
@@
    The Lagranzian $\mathbb{L}_{\sigma}$ is biconic along $\partial_{\infty}\mathbb{L}_{\sigma}$.
@@
    The FLTZ Lagranzian $\mathbb{L}(\Sigma)$ is biconic along each $\partial_{\infty}\mathbb{L}_{\sigma}$.
@@
    For each $\sigma\in\Sigma$ the local coordinates $\eta_{\sigma}$ define a Weinstein hypersuface
@@
Remark 3.12.
As explained above, we glue two Weinstein domains along the standard contact neighborhoods of their Weinstein hypersurfaces. Here, each $R_{\sigma}$ is a Weinstein thickening of $\partial_{\infty}\mathbb{L}_{\sigma}$. Lemma 3.11(4) guarantees compatibility of local Weinstein pairs $(T^{*}\widehat{M},R_{\sigma})$ with the union $\mathbb{L}(\Sigma)=\bigcup_{\sigma\in\Sigma}\mathbb{L}_{\sigma}$.
@@ 4.1. Sectorial covers
Due to [GPS2], given a Liouville sector $X$, if it admits a Weinstein sectorial cover, then the wrapped Fukaya category $\mathcal{W}(X)$ can be computed by gluing that of pieces in the cover.
Definition 4.1 ([GPS2, Definition 12.2]).
Let $X$ be a Liouville manifold-with-boundary. A collection of cylindrical hypersurfaces $H_{1},\ldots,H_{n}\subset X$ is sectorial if their characteristic foliations $C_{1},\ldots,C_{n}$ are $\omega$-orthogonal over their intersections and there exist functions $I_{i}\colon\operatorname{Nbd}^{Z}_{X}(H_{i})\to\mathbb{R}$ linear near infinity satisfying
(4.1) $dI_{i}|_{C_{i}}\neq 0,\ dI_{i}|_{C_{j}}=0\ \text{for}\ i\neq j,\ \{I_{i},I_{j}\}=0.$
We also allow immersed hypersurfaces $H\to X$, with $I_{i}$ defined on $\operatorname{Nbd}^{Z}_{X}(H_{i})$ regarded as an immersed codimension zero submanifold of $X$, and the subscripts $i$ in the identities (4.1) indexing the “local branches” of $H$.
Definition 4.2 ([GPS2, Definition 12.14]).
A Liouville sector-with-sectorial-corners is a Liouville manifold-with-corners whose boundary, regarded as an immersed hypersurface, is sectorial.
@@
Let $X$ be a Liouville manifold-with-boundary. Suppose that $X$ admits a cover $X=\bigcup^{n}_{i=1}X_{i}$ by Liouville sectors-with-sectorial-corners $X_{i}\subset X$, with precisely two faces $\partial^{1}X_{i}=X_{i}\cap\partial X$ and the point set topological boundary $\partial^{2}X_{i}$ of $X_{i}$ meeting along the corner locus $\partial X\cap\partial^{2}X_{i}=\partial^{1}X_{i}\cap\partial^{2}X_{i}$. Such a cover is sectorial if the collection of boundaries $\partial X,\partial^{2}X_{1},\ldots,\partial^{2}X_{n}$ is sectorial. We also allow $X_{i}$ without corners, requiring $\partial^{2}X_{i}$ to be disjoint from $\partial X$, $\partial X,\partial^{2}X_{1},\ldots,\partial^{2}X_{n}$ to be sectorial and $\bigcup^{n}_{i=1}X_{i}\hookrightarrow X$ to be a trivial inclusion.
@@
ranging over $J\sqcup K\sqcup L=\{1,\ldots,n\}$. The closure of each $X_{J,K,L}$ is a submanifold-with-corners, whose symplectic reduction is a Liouville sector-with-sectorial-corners.
Definition 4.4 ([GPS2, Definition 12.15]).
A sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$ is Weinstein if the convex completions of all of the symplectic reductions of strata (4.2) are Weinstein up to deformation.
Theorem 4.5 ([GPS2, Theorem 1.35]).
For any Weinstein sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$ of a Weinstein sector $X$, the induced functor
@@
is a pretriangulated equivalence. Moreover, the same holds for any mostly Legendrian stop $\mathfrak{r}\subset(\partial_{\infty}X)^{\circ}$ which is disjoint from each $\partial X_{i}$ and from a neighborhood of $\partial X$.
@@ 4.2. Alternative description of $\widetilde{\mathbf{W}}(\Phi)$ and $\bar{\pi}$
Recall that $\widetilde{\mathbf{W}}(\Phi)$ is a Weinstein manifold-with-boundary obtained by Weinstein handle attachments, special cases of gluing of Weinstein pairs. The boundary comes from conic completion of the boundaries of the contact surroundings of $T^{*}\mathcal{L}_{k}$ in $\partial_{\infty}{\mathbf{W}}(\Phi_{k-1})$. Below, preserving a sufficiently small neighborhood of $\widetilde{\mathbb{L}}(\Phi)$, we will modify the construction in order to obtain a Weinstein manifold without boundary. See [Eli, Figure 3.1] to get the picture of our modification. As in the original construction, we proceed by induction on $k$.
@@ 4.3. Construction
We construct a Weinstein sectorial cover of the modification $\widetilde{\mathbf{W}}(\Phi)$ as the inverse image under $\bar{\pi}$ of a certain cover of $\Phi$. For expository reason, we will assume that $\Phi$ is closed. Since each stratum is closed, one can take its barycenter. Connect by an edge inside $\Phi$ the barycenter on each top dimensional stratum with that on $1$-dimensional lower strata adjacent to it. Repeat the process inductively until reaching the barycenters of $1$-strata. There might be some lower dimensional closed strata which are not adjacent to higher dimensional strata. Beginning the process also from such strata, we obtain a partition of $\Phi$ defined by the additional edges connecting the barycenters and the barycenters on those $1$-strata which are not adjacent to higher dimensional strata. The partition divides $\Phi$ into some finite number of connected components.
For each 00-stratum $P$, there is a unique connected component containing $P$. Take a suitable slight enlargement of the rounding the corners of its closure. We denote by $\widetilde{\mathbf{W}}(P)$ and $\partial^{2}\widetilde{\mathbf{W}}(P)$ the inverse image under $\bar{\pi}$ respectively of the slight enlargement and of its boundary. Perturbing these boundaries if necessary, we may assume that $\partial^{2}\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ are disjoint or intersect transversely for any pair $P,P^{\prime}$ of 00-strata. Perturb each edge connecting the barycenter of a top dimensional stratum $S^{(n)}$ to that of a $1$-dimensional lower stratum $S^{(n-1)}$ so that it becomes locally in $\Phi_{n-1}$ perpendicular to $S^{(n-1)}$. More precisely, the edge becomes parallel to the normal direction in the handle $\Sigma_{S^{(n-1)}}\times S^{(n-1)}\subset\Phi_{n-1}\setminus\Phi_{n-2}$ to $S^{(n-1)}$. This is always possible since $\Phi$ is smoothly normally conical. Repeat the process inductively until reaching the barycenters of $1$-strata in $\Phi_{1}$. Perform the same perturbation to each edge starting from the barycenter of a lower dimensional closed stratum which is not adjacent to higher dimensional strata. We use the same symbols $\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P)$ to denote the results.
@@
We adapt the idea in [GPS2, Example 1.33]. Recall that we are assuming $\Phi$ to be closed. For simplicity, we will further assume that there is no lower dimensional strata which are not adjacent to higher dimensional strata. Since by construction $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ is away from $P$, we may assume that $\partial^{2}\widetilde{\mathbf{W}}(P)$ is disjoint from $T^{*}\widehat{M}_{P}\subset\widetilde{\mathbf{W}}(P)\cap\widetilde{\mathbf{W}}(\Phi_{0})$. Let $I$ be a $1$-stratum which intersects $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ at a point $p$. Due to the above perturbation, locally in $\Phi_{1}$ the image $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ becomes perpendicular to $I$. Since $\bar{\pi}$ kills the base direction of $T^{*}\widehat{M}_{I}$ and the fiber direction of $T^{*}I_{\circ}$, over the neighborhood of $I$ the boundary $\partial^{2}\widetilde{\mathbf{W}}(P)$ is isomorphic to a product of $T^{*}\widehat{M}_{I}$ and the fiber $T^{*}_{p}I_{\circ}$. Inductively, let $S^{(k)}$ be a $k$-stratum adjacent to a $(k-1)$-stratum $S^{(k-1)}$ which intersects $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$. Due to the above perturbation, locally in $\Phi_{k}$ the image $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ becomes perpendicular to $S^{(k)}$. Since $\bar{\pi}$ kills the base direction of $T^{*}\widehat{M}_{S^{(k)}}$ and the fiber direction of $T^{*}S^{(k)}_{\circ}$, over the neighborhood of $S^{(k)}$ the boundary $\partial^{2}\widetilde{\mathbf{W}}(P)$ is isomorphic to a product of $T^{*}\widehat{M}_{S^{(k)}}$ and the restriction $T^{*}S^{(k)}_{\circ}|_{\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))\cap S^{(k)}}$. By construction locally in $\Phi_{(k-1)}$ the intersection $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))\cap S^{(k)}$ becomes perpendicular to $S^{(k-1)}$, the handle attachment $\mathbf{W}(\Phi_{k-1})\#_{\mathcal{L}_{k-1}}(T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ})$ glues this product to the gluing of the products in the previous steps.
@@
Rewinding the process, one sees that $\partial^{2}\widetilde{\mathbf{W}}(P)$ is obtained by extending via the handle attachments the union of the inverse images under $\bar{\pi}$ of hypersurfaces in the intersection of $\Phi_{n}\setminus\Phi_{n-1}$ and the connected component containing $P$. We claim that this hypersurface $\partial^{2}\widetilde{\mathbf{W}}(P)$ is sectorial. Since the newly attached pieces are cylindrical, so is $\partial^{2}\widetilde{\mathbf{W}}(P)$. It remains to show the existence of an $\alpha$-defining function on $\partial^{2}\widetilde{\mathbf{W}}(P)$. By [GPS2, Construction 12.17] we may assume $\partial^{2}\widetilde{\mathbf{W}}(P)$ to be one of the hypersurfaces-with-corners dividing $\widetilde{\mathbf{W}}(\Phi)$ into the original connected components. Let $I$ be a $1$-stratum adjacent to $P$ and $H_{P,I}$ any of such hypersurfaces intersecting $I$. Fix the standard coordinates $(x,y)$ on the symplectic manifold $T^{*}I_{\circ}$. Unwinding the filtration of $\Phi$ and the construction of $\widetilde{\mathbf{W}}(\Phi)$, one sees that $H_{P,I}$ is given by $\{x=0\}$ up to perturbation. We denote by $y_{P,I}$ the function on a neighborhood of $H_{P,I}$ which is canonically induced by $y\colon T^{*}I_{\circ}\to\mathbb{R}$. Clearly, $y_{P,I}$ is linear near infinity. The Hamiltonian vector field $X_{y_{P,I}}$ is outward pointing along $H_{P,I}$, as it is given by a lift of $\partial_{x}$ on $T^{*}I_{\circ}$. Hence $H_{P,I}$ is sectorial. Thus the Weinstein manifold-with-boundary $\widetilde{\mathbf{W}}(P)$ satisfies the condition in Definition 2.1 to be a Liouville sector. ∎
@@
Suppose that there is a lower dimensional closed stratum which is not adjacent to higher dimensional strata. Let $S^{(k)}$ be such a stratum. Then for each 00-stratum $P$ adjacent to it, $\partial^{2}\widetilde{\mathbf{W}}(P)$ gets contribution from the extensions via the handle attachments of the inverse images under $\bar{\pi}$ of hypersurfaces in $S^{(k)}$.

### v3

@@ 2.1. Fanifolds
    $\Phi$ is conical in the complement of a compact subset [GS23, Page 37].
@@
One can identify the exit path category $\operatorname{Exit}(\Sigma)$ with $\operatorname{Fan}^{\twoheadrightarrow}_{\Sigma/}$ as a poset via $\sigma\mapsto[\Sigma\mapsto\Sigma/\sigma]$. In addition, the normal geometry to $\sigma$ is the geometry of $\Sigma/\sigma$. Due to the condition (v), $\operatorname{Exit}_{S}(\Phi)$ is equivalent to the poset $\operatorname{Exit}(\Sigma_{S})$, the full subcategory of exit paths starting at $S$ contained inside a sufficiently small neighborhood of $S$.
@@
Suppose that $\Phi_{k-1}$ has the desired filtration $\Phi_{0}\subset\Phi_{1}\subset\cdots\subset\Phi_{k-1}$, where $\Phi_{0}$ is the disjoint union $\bigsqcup_{P}\Sigma_{P}$ for all 0-strata $P\subset\Phi$ equipped with the canonically induced fanifold structure. The ideal boundary $\partial_{\infty}S$ might have some subset $\partial_{in}S$ which is in the direction to the interior of $\Phi$. Perform the gluing $\Phi_{k-1}\#_{(\Sigma_{S}\times\partial_{in}S)}(\Sigma_{S}\times S)$ which equals $\Sigma_{S}\times S$ when $\Phi_{k-1}$ is empty and equals $\Phi_{k-1}$ unless $S$ is an interior $k$-stratum, i.e. $\partial_{in}S=\partial S_{\circ}$ where $S_{\circ}=S\setminus\Phi_{k-1}$ is a manifold-with-boundary. Let
@@ 2.2. Weinstein handle attachments
Throughout the paper, we work with Liouville manifolds of finite type so that any Liouville manifold $X$ will be the completion of some Liouville domain $X_{0}$. Then its skeleton $\operatorname{Core}(X)$ will always be compact. We use the symbol $\partial_{\infty}X$ to denote both the ideal boundary of $X$ and the actual boundary of $X_{0}$ to preserve the symbols $\partial X_{0},\partial X$. All Liouville manifolds of our interest are Weinstein. Namely, the Liouville vector field $Z$ of each $X$ is gradient-like for a Morse–Bott function $\phi\colon X\to\mathbb{R}$ which is constant on the cylindrical end $(\mathbb{R}_{t\geq 0}\times\partial_{\infty}X,e^{t}(\lambda|_{\partial_{\infty}X}))$. Hence $\operatorname{Core}(X)$ will also be isotropic by [CE12, Lemma 11.13(a)].
@@
where the Liouville flow on $W$ gets identified with the translation action on $\mathbb{R}_{\leq 0}$. By [Eli, Proposition 2.9] one can modify the Weinstein structure near $\mathcal{L}$ so that the Liouville flow gets identified with the cotangent scaling on $T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$. We denote by $\widetilde{W}$ its conic completion, which is a Weinstein sector by [GPS1, Lemma 2.13]. Then $\eta$ induces a neighborhood $\widetilde{\eta}\colon\operatorname{Nbd}_{\partial_{\infty}\widetilde{W}}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}$ which extends to a neighborhood $\widetilde{\xi}\colon\operatorname{Nbd}_{\widetilde{W}}(\mathcal{L})\hookrightarrow J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}$.
@@
where $\operatorname{Core}(W,\mathcal{L})$ denotes the relative skeleton of the stopped Weinstein domain $(W,\mathcal{L})$, i.e. the disjoint union $\operatorname{Core}(W)\sqcup\mathbb{R}\mathcal{L}\subset W$ of $\operatorname{Core}(W)$ and the saturation of $\mathcal{L}$ by the Liouville flow.
@@
Let $W$ be a Weinstein domain with a smooth Legendrian $\mathcal{L}\subset\partial_{\infty}W$. Then a Lagrangian $\mathbb{L}\subset W$ is biconic along $\mathcal{L}$ if and only if it is conic and there exists some $\xi\colon\operatorname{Nbd}_{W}(\mathcal{L})\hookrightarrow T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0})$ such that
$\xi(\mathbb{L}\cap\operatorname{Nbd}_{W}(\mathcal{L}))\subset T^{*}\mathcal{L}\times\{0\}\times\mathbb{R}_{\leq 0}\subset J^{1}\mathcal{L}\times\mathbb{R}_{\leq 0}\cong T^{*}(\mathcal{L}\times\mathbb{R}_{\leq 0}),$
where the Liouville flow on $W$ gets identified with the translation action on $\mathbb{R}_{\leq 0}$.
@@
For a closed manifold $\widehat{M}$ and a compact manifold-with-boundary $S$, consider the Weinstein manifold-with-boundary $\widetilde{W^{\prime}}=T^{*}\widehat{M}\times T^{*}S$ with a smooth Legendrian $\mathcal{L}=\widehat{M}\times\partial S$ taken to be a subset of the zero section. Here, we equip $T^{*}S$ the tautological Liouville structure so that the Liouville vector field is the generator of fiberwise radial dilation. Then the Liouville flow on $\widetilde{W^{\prime}}$ near $\mathcal{L}$ is the cotangent scaling. We write $W^{\prime}$ for a domain completing to $\widetilde{W^{\prime}}$. One can check that the above gluing procedure carries over, although $\mathcal{L}$ does not belong to $\partial_{\infty}W^{\prime}$.
@@
Let $\Sigma\subset M_{\mathbb{R}}$ be a rational polyhedral fan and $\widehat{M}$ the real $n$-torus $\operatorname{Hom}(M,\mathbb{R}/\mathbb{Z})=M^{\vee}_{\mathbb{R}}/M^{\vee}$. Recall from [FLTZ12, Section 3.1] that the FLTZ Lagrangian is the union
@@
    The Lagrangian $\mathbb{L}_{\sigma}$ is biconic along $\partial_{\infty}\mathbb{L}_{\sigma}$.
@@
    The FLTZ Lagrangian $\mathbb{L}(\Sigma)$ is biconic along each $\partial_{\infty}\mathbb{L}_{\sigma}$.
@@
    For each $\sigma\in\Sigma$ the local coordinates $\eta_{\sigma}$ define a Weinstein hypersurface
@@ 3.1. Sectorial covers
Throughout the paper, for a stopped Liouville sector $(X,\mathfrak{f})$ we denote by $\mathcal{W}(X,\mathfrak{f})$ the partially wrapped Fukaya category [GPS2, definition 2.31]. When $\mathfrak{f}=\emptyset$, it gives the wrapped Fukaya category $\mathcal{W}(X)$ of a Liouville sector $X$ [GPS1, Definition 3.36]. Due to [GPS2], given a Liouville sector $X$, if it admits a Weinstein sectorial cover, then the wrapped Fukaya category $\mathcal{W}(X)$ can be computed by gluing that of pieces in the cover.
@@
Let $X$ be a Liouville manifold-with-boundary. Suppose that $X$ admits a cover $X=\bigcup^{n}_{i=1}X_{i}$ by Liouville sectors-with-sectorial-corners $X_{i}\subset X$ [GPS2, Definition 12.14], with precisely two faces $\partial^{1}X_{i}=X_{i}\cap\partial X$ and the point set topological boundary $\partial^{2}X_{i}$ of $X_{i}$ meeting along the corner locus $\partial X\cap\partial^{2}X_{i}=\partial^{1}X_{i}\cap\partial^{2}X_{i}$. Such a cover is sectorial if the collection of boundaries $\partial X,\partial^{2}X_{1},\ldots,\partial^{2}X_{n}$ is sectorial in the sense of [GPS2, Definition 12.2]. We also allow $X_{i}$ without corners, requiring $\partial^{2}X_{i}$ to be disjoint from $\partial X$ and $\partial X,\partial^{2}X_{1},\ldots,\partial^{2}X_{n}$ to be sectorial and $\bigcup^{n}_{i=1}X_{i}\hookrightarrow X$ to be a trivial inclusion.
@@
ranging over $J\sqcup K\sqcup L=\{1,\ldots,n\}$. The closure of each $X_{J,K,L}$ is a submanifold-with-corners, whose symplectic reduction is a Liouville sector-with-sectorial-corners. Recall that a sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$ is Weinstein [GPS2, Definition 12.15] if the convex completions of all of the symplectic reductions of strata (3.1) are Weinstein up to deformation. For any Weinstein sectorial cover $X=\bigcup^{n}_{i=1}X_{i}$ of a Weinstein sector $X$, the induced functor
@@
is a pretriangulated equivalence [GPS2, Theorem 1.35].
@@ 3.2. Alternative description of $\widetilde{\mathbf{W}}(\Phi)$ and $\bar{\pi}$
Recall that $\widetilde{\mathbf{W}}(\Phi)$ is a stopped Weinstein sector obtained by Weinstein handle attachments. See also Remark 2.6. The boundary comes from conic completion of the boundaries of the contact surroundings of $T^{*}\mathcal{L}_{k}$ in $\partial_{\infty}{\mathbf{W}}(\Phi_{k-1})$. Below, preserving a sufficiently small neighborhood of $\widetilde{\mathbb{L}}(\Phi)$, we will modify the construction in order to obtain a stopped Weinstein manifold. See [Eli, Figure 3.1] to get the picture of our modification. As in the original construction, we proceed by induction on $k$.
@@ 3.3. Construction
We construct a Weinstein sectorial cover of the modification $\widetilde{\mathbf{W}}(\Phi)$ as the inverse image under $\bar{\pi}$ of a certain cover of $\Phi$. For expository reasons, we will assume that $\Phi$ is closed. Since each stratum is closed, one can take its barycenter. Connect by an edge inside $\Phi$ the barycenter on each top dimensional stratum with that on $1$-dimensional lower strata adjacent to it. Repeat the process inductively until reaching the barycenters of $1$-strata. There might be some lower dimensional closed strata which are not adjacent to higher dimensional strata. Beginning the process also from such strata, we obtain a partition of $\Phi$ defined by the additional edges connecting the barycenters and the barycenters on those $1$-strata which are not adjacent to higher dimensional strata. The partition divides $\Phi$ into some finite number of connected components.
For each 0-stratum $P$, there is a unique connected component containing $P$. Take a suitable slight enlargement of rounding of corners of its closure. We denote by $\widetilde{\mathbf{W}}(P)$ and $\partial^{2}\widetilde{\mathbf{W}}(P)$ the inverse image under $\bar{\pi}$ respectively of the slight enlargement and of its boundary. Perturbing these boundaries if necessary, we may assume that $\partial^{2}\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P^{\prime})$ are disjoint or intersect transversely for any pair $P,P^{\prime}$ of 0-strata. Perturb each edge connecting the barycenter of a top dimensional stratum $S^{(n)}$ to that of a $1$-dimensional lower stratum $S^{(n-1)}$ so that it becomes locally in $\Phi_{n-1}$ perpendicular to $S^{(n-1)}$. More precisely, the edge becomes parallel to the normal direction in the handle $\Sigma_{S^{(n-1)}}\times S^{(n-1)}\subset\Phi_{n-1}\setminus\Phi_{n-2}$ to $S^{(n-1)}$. This is always possible since $\Phi$ is smoothly normally conical. Repeat the process inductively until reaching the barycenters of $1$-strata in $\Phi_{1}$. Perform the same perturbation to each edge starting from the barycenter of a lower dimensional closed stratum which is not adjacent to higher dimensional strata. We use the same symbols $\widetilde{\mathbf{W}}(P),\partial^{2}\widetilde{\mathbf{W}}(P)$ to denote the results.
@@
We adapt the idea in [GPS2, Example 1.33]. Recall that we are assuming $\Phi$ to be closed. For simplicity, we will further assume that there is no lower dimensional strata which are not adjacent to higher dimensional strata. Since by construction $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ is away from $P$, we may assume that $\partial^{2}\widetilde{\mathbf{W}}(P)$ is disjoint from $T^{*}\widehat{M}_{P}\subset\widetilde{\mathbf{W}}(P)\cap\widetilde{\mathbf{W}}(\Phi_{0})$. Let $I$ be a $1$-stratum which intersects $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ at a point $p$. Due to the above perturbation, locally in $\Phi_{1}$ the image $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ becomes perpendicular to $I$. Since $\bar{\pi}$ kills the base direction of $T^{*}\widehat{M}_{I}$ and the fiber direction of $T^{*}I_{\circ}$, over the neighborhood of $I$ the boundary $\partial^{2}\widetilde{\mathbf{W}}(P)$ is isomorphic to a product of $T^{*}\widehat{M}_{I}$ and the fiber $T^{*}_{p}I_{\circ}$. Inductively, let $S^{(k)}$ be a $k$-stratum adjacent to a (k1)(k-1)-stratum $S^{(k-1)}$ which intersects $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$. Due to the above perturbation, locally in $\Phi_{k}$ the image $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))$ becomes perpendicular to $S^{(k)}$. Since $\bar{\pi}$ kills the base direction of $T^{*}\widehat{M}_{S^{(k)}}$ and the fiber direction of $T^{*}S^{(k)}_{\circ}$, over the neighborhood of $S^{(k)}$ the boundary $\partial^{2}\widetilde{\mathbf{W}}(P)$ is isomorphic to a product of $T^{*}\widehat{M}_{S^{(k)}}$ and the restriction $T^{*}S^{(k)}_{\circ}|_{\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))\cap S^{(k)}}$. By construction locally in $\Phi_{(k-1)}$ the intersection $\bar{\pi}(\partial^{2}\widetilde{\mathbf{W}}(P))\cap S^{(k)}$ becomes perpendicular to $S^{(k-1)}$, the handle attachment $\mathbf{W}(\Phi_{k-1})\#_{\mathcal{L}_{k-1}}(T^{*}\widehat{M}_{S^{(k)}}\times T^{*}S^{(k)}_{\circ})$ glues this product to the gluing of the products in the previous steps.
@@
Rewinding the process, one sees that $\partial^{2}\widetilde{\mathbf{W}}(P)$ is obtained by extending via the handle attachments the union of the inverse images under $\bar{\pi}$ of hypersurfaces in the intersection of $\Phi_{n}\setminus\Phi_{n-1}$ and the connected component containing $P$. We claim that this hypersurface $\partial^{2}\widetilde{\mathbf{W}}(P)$ is sectorial. Since the newly attached pieces are cylindrical, so is $\partial^{2}\widetilde{\mathbf{W}}(P)$. It remains to show the existence of an $\alpha$-defining function on $\partial^{2}\widetilde{\mathbf{W}}(P)$. By [GPS2, Construction 12.17] we may assume $\partial^{2}\widetilde{\mathbf{W}}(P)$ to be one of the hypersurfaces-with-corners dividing $\widetilde{\mathbf{W}}(\Phi)$ into the original connected components. Let $I$ be a $1$-stratum adjacent to $P$ and $H_{P,I}$ any of such hypersurfaces intersecting $I$. Fix the standard coordinates $(x,y)$ on the symplectic manifold $T^{*}I_{\circ}$. Unwinding the filtration of $\Phi$ and the construction of $\widetilde{\mathbf{W}}(\Phi)$, one sees that $H_{P,I}$ is given by $\{x=0\}$ up to perturbation. We denote by $y_{P,I}$ the function on a neighborhood of $H_{P,I}$ which is canonically induced by $y\colon T^{*}I_{\circ}\to\mathbb{R}$. Clearly, $y_{P,I}$ is linear near infinity. The Hamiltonian vector field $X_{y_{P,I}}$ is outward pointing along $H_{P,I}$, as it is given by a lift of $\partial_{x}$ on $T^{*}I_{\circ}$. Hence $H_{P,I}$ is sectorial. Thus the Weinstein manifold-with-boundary $\widetilde{\mathbf{W}}(P)$ satisfies the condition in [GPS1, Definition 2.4] to be a Liouville sector. ∎
@@
Suppose that there is a lower dimensional closed stratum which is not adjacent to higher dimensional strata. Let $S^{(k)}$ be such a stratum. Then for each 0-stratum $P$ adjacent to it, $\partial^{2}\widetilde{\mathbf{W}}(P)$ gets contribution from the extensions via the handle attachments of the inverse images under $\bar{\pi}$ of hypersurfaces in $S^{(k)}$.